OpenStudy (anonymous):

Complex Variables Find where in the complex plane the following function is analytic:

OpenStudy (anonymous):

\[\frac{ 1 }{ (y - ix + 1 + i + z)^4 }\] i have the textbook's answer if needed but I don't understand why. please explain ^_^

OpenStudy (superdavesuper):

Ur question is a lil unclear - which IS the function and what is the definition of "analytic"? I assume z is the function....which is confusing as it could mean the z-axis in a 3-D coordinate system. Please clarify.

OpenStudy (anonymous):

it is a function f(z) where z is a function of x and y

OpenStudy (superdavesuper):

yup so z=f(x,y) I presume analytic means that the function is not infinity. So the region is wherever on the complex plane except when z = -y + ix -1 -i because it will become 1/0.

OpenStudy (anonymous):

that isn't the textbook answer though..

OpenStudy (amoodarya):

did z=x+iy ?

OpenStudy (amoodarya):

if yes put z=x+iy then simplify (x+y+1) +i(-x+1+y) =0+0i so solve the system of equation you will find x=0 , y=-1 so your function is analytic always but not in z=-i

OpenStudy (anonymous):

yes z = x + iy thank you very much ^_^

OpenStudy (amoodarya):

do you understand ?